OpenStudy (anonymous):

I need help solving: If sin(theta)=12/13, cos >0 then what would cos(2theta) be?

OpenStudy (anonymous):

\[\sin(2\theta)=1-2\sin^2(x)\]

OpenStudy (anonymous):

replace \(\sin(x)\) by \(\frac{12}{13}\) and do the arithmetic

OpenStudy (anonymous):

ok that was not what i meant to write i meant \[\cos(2\theta)=1-2\sin^2(x)\]

OpenStudy (anonymous):

your only job now is to compute \[1-\left(\frac{12}{13}\right)^2\]

OpenStudy (anonymous):

that business about what quadrant you are in is a red herring